TRANSIENT FREQUENCY ABSORPTION STUDIES OF MOLECULES PRODUCED IN A LASER ABLA-TION - SUPERSONIC EXPANSION SOURCE

\underline{T. C. STEIMLE}, Department of Chemistry and Biochemistry, Arizona State University, Tempe, AZ, 85287-1604; M. COSTEN, G. E. HALL and T. J. SEARS, Department of Chemistry Brookhaven National Laboratory, Upton, NY 11973.

The supersonic free jet expansion of a laser ablation/reaction source has been coupled to sensitive transient frequency modulation absorption spectroscopy^{1,2} for the first time. The (0,0)A^{1}Σ − X^{1}Σ band system (ν= 12643 cm^{−1}) of PtC and the (0,0) C^{3}Δ − D^{3}Δ band system of TiS (ν= 12065 cm^{−1}) were recorded. Excellent S/N could be achieved using a single pass through the expansion 5 cm from the ablation source. Narrow spectral features(< 200 MHz) were achieved by appropriate temporal gating of the transient signal. The observed absorption lineshapes exhibit a dependence on the intrinsic transition moment and therefore the probe laser power, which facilitated the assignment of the optical spectrum.

1 J.C. Bloch, R.W. Field, G.E. Hall and T.J. Sears, J. Chem. Phys. 101, 1717 (1994). 2 A.J. Marr, T.J. Sears, B-C Chang, J. Chem. Phys. 109, 3431 (1998).

Acknowledgement: This work was carried out at Brookhaven National Laboratory under contract DE-AC02-98CH10886 with the U.S. Department of Energy and supported by its Division of Chemical Sciences, Office of Basic Energy Sciences.